OpenStudy (nannie):

Simplify y^2-5y+6/y^+2y-15 PLEASE!

OpenStudy (anonymous):

=(y-2)(y-3)/(y-3)(y+5) = (y-2)/(y+5)

OpenStudy (anonymous):

Y^2 - 5y + 6 = (y -2)(y-3) y^2 + 2x - 15 = (y - 3)(y + 5)\[\ \frac{(y-2)(y-3)}{(y-3)(y+5)} = \ \frac{y-2}{y+5} \]

OpenStudy (dboyette):

(y^(3)-3y^(2)-15y+6)/(y)

OpenStudy (nannie):

Thank you both! I had it reversed. Thanks again

OpenStudy (anonymous):

you're welcome!